OpenStudy (anonymous):

decide whether y is a function of x. y=-3x-4

OpenStudy (anonymous):

Solve for x. y+4=-3x y+4/-3=x So, x=(y+4)/-3 Is that a function?

OpenStudy (anonymous):

It may be a little out of context, but If im interpreting this right, then by the definition of a function, if there is no more than 1value of y that can equal 1 or more x value, then y is a function of x. for example. y^2 = x is not a function of x because: y^2 = x --> y =( + or -)sqrt(x). that means there is more than 1 possible value of y for a single value of x. the same idea applies to finding if x is a function of y. if there can be only one x value or 1 or more values of y, then x can also be a function of x.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2nd and 4th quadrant of x and y has a function